# Comparing Bad Apples to Good Oranges: Aligning Large Language Models via Joint Preference Optimization

### v1

@@ Appendix E Dataset Statistics
Appendix F ChatGPT Prompts
You are a helpful following assistant whose goal is to select the preferred output for a given instruction. Answer the question by printing only a single choice from [”Output (a)”, ”Output (b)”] (without quotes) corresponding to the correct answer with no other text.
#### Annotation Guideline
You will read examples, which are composed of the following:
1. an Instruction we give to the AI system
2. Output (a), the first output from the AI system
3. Output (b), the second output from the AI system
### Instruction:
${instruction}$
### Output (a):
${output_{1}}$
### Output (b):
${output_{2}}$
## Preferred Output:
You are a helpful following assistant whose goal is to select the preferred output for a given instruction. Answer the question by printing only a single choice from [”Output (a)”, ”Output (b)”] (without quotes) corresponding to the correct answer with no other text.
#### Annotation Guideline
You will read examples, which are composed of the following:
### Instruction (a):
${instruction_{1}}$
### Output (a):
${output_{1}}$
### Instruction (b):
${instruction_{2}}$
### Output (b):
${output_{2}}$
## Preferred Output:
Appendix G Human Annotation Platform
We present the screenshots for the human interface in the Figure 6 (conditional rankings) and Figure 7 (joint ranking preferences over instruction-response pairs).
Figure 6: Human annotation interface for Conditional Rankings
Appendix H Qualitative Examples
@@
H.1 Anthropic-Helpful Examples
We present the qualitative examples for the preferences acquired for the Anthropic-helpful dataset in Figure 8, 9, and 10. We present our observations in the figure captions.
Figure 8: Interplay between the conditional rankings and joint rankings and reasoning acquired from the human annotators for the Anthropic-Helpful dataset. In this example, we find that the response B and D are rejected under the conditional rankings. When asked to compare the response B and D, humans consider that the response B answers Instruction 1 better than response D answers Instruction 2. This indicates that the joint preference humans elicits a decisive feedback between two responses that were rejected under the conditional rankings.
@@ H.1 Anthropic-Helpful Examples
H.2 TL;DR Summarization Examples
We present the qualitative examples for the preferences acquired for the TL;DR summarization dataset in Figure 11, 12, and 13. We present our observations in the figure captions.
@@ H.2 TL;DR Summarization Examples
Appendix I Alignment Training Details
I.1 Supervised Finetuning Details
We present the SFT details in  Table 4. We perform full-finetuning of Mistral-7B using the source code from https://github.com/abacaj/fine-tune-mistral.
@@ I.1 Supervised Finetuning Details
I.2 Dove
We present the training details for Dove preference optimization objective in the Table 5. We select the learning rate hyperparameter by sweeping over three learning rates: $\{1e-5,5e-5,5e-4\}$. We utilize the TRL library [52] for the DPO source code.
@@ I.2 Dove
Optimizer AdamW [27]
@@
Table 5: Training details for Dove preference optimization objective. We use the identical settings for DPO.
Appendix J Trends with Data Scaling
We aim to understand the impact of increasing the number of preferences collected jointly over instruction-response pairs, for non-identical instructions, on the win-rate against the reference summaries in the TL;DR summarization dataset using Dove algorithm. We present the results in Figure 14 for the sampling temperature of $0.001$. We find that the win-rate scales from $42.4\%$ to $71.7\%$ as the size of the dataset increases from $100$ to $9000$ comparisons. We also observe that the change in the win-rate is within $1\%$ when the dataset size increases from $4000$ to $9000$. This highlights that the performance gains are non-linear with the dataset size. In the future, it would be pertinent to explore techniques for selecting a subset of joint preference comparisons that result in maximum performance gains.
Figure 14: Results for scaling the feedback data size on TL;DR summarization dataset. We find that the win-rate improves with the increase in the dataset size using the Dove preference optimization objective.

### v2

@@ Appendix C Dataset Statistics
Appendix D Proof for JPO subsuming DPO
We highlight a result that reduces JPO into DPO when the prompts are the same in Lemma E.1.
Lemma D.1.
Under the case where $\mathcal{D}_{X}=\{(I_{i},R_{i},I_{i},R_{j})\}$, that is, prompts are the same for preferred and not-preferred prompt generation pairs, $\mathcal{L}_{\textsc{DPO}}(\theta;\mathcal{D}_{C},\beta,p_{\text{ref}})=\mathcal{L}_{\textsc{JPO}}(\theta;\mathcal{D}_{X},\beta,p_{\text{ref}})$, where $\mathcal{D}_{C}=\{(I_{j},R_{j}^{w},R_{j}^{\ell})\}$.
Proof.
$\mathcal{L}_{\textsc{JPO}}(\theta;\mathcal{D}_{X},\beta,p_{\text{ref}})$ $=\mathbb{E}_{(I_{j}^{w},R_{j}^{w},I_{j}^{\ell},R_{j}^{\ell})\sim\mathcal{D}_{X}}\left[\log\left(\sigma\left(\beta\log\frac{p_{\theta}(R_{i}^{w},I_{i}^{w})}{p_{\text{ref}}(R_{i}^{w},I_{i}^{w})}-\beta\log\frac{p_{\theta}(R_{j}^{\ell},I_{j}^{\ell})}{p_{\text{ref}}(R_{j}^{\ell},I_{j}^{\ell})}\right)\right)\right]$ (3)
$=\mathbb{E}_{(I_{j}^{w}R_{j}^{w},I_{j}^{\ell},R_{j}^{\ell})\sim\mathcal{D}_{X}}\left[\log\left(\sigma\left(\beta\log\frac{p_{\theta}(R_{i}^{w}|I_{i}^{w})p_{\theta}(I_{i}^{w})}{p_{\text{ref}}(R_{i}^{w}|I_{i}^{w})p_{\text{ref}}(I_{i}^{w})}-\beta\log\frac{p_{\theta}(R_{j}^{\ell}|I_{j}^{\ell})p_{\theta}(I_{j}^{\ell})}{p_{\text{ref}}(R_{j}^{\ell},I_{j}^{\ell})p_{\text{ref}}(I_{j}^{\ell})}\right)\right)\right]$ (4)
Appendix E JPO on AlpacaEval2 Leaderboard
We train Mistral-7B base model on the UltraChat-200K dataset Ding et al. (2023) to get the SFT (reference) model. Subsequently, we utilize the conditional preference dataset, Ultrafeedback-binarized (60K instances) Cui et al. (2023) to align the SFT model using DPO as the baseline algorithm. Specifically, we utilize the training setup highlighted in the alignment handbook for SFT and DPO Tunstall et al. (2023). Since JPO algorithm allows access to joint preferences, we construct non-identical instruction-response tuples by pairing a chosen instruction-response ($I_{chosen},R_{chosen}$) with a rejected instruction-response ($I_{reject},R_{reject}$) from the Ultrafeedback dataset. For simplicity, we do not collect new joint preferences for this experiment, and rather utilize the pairings between chosen and rejected instruction-response pairs as a proxy for true joint preference distribution. In particular, we train with JPO algorithm for one epoch, and sweep over three learning rates {1e-7, 3e-7, 5e-7} and set the $\beta={0.01}$. Post-training, we sample responses from the SFT model, DPO-aligned LLM, KTO-aligned LLM, and JPO-aligned LLM for the instructions in the AlpacaEval2 with a temperature of $0.7$.
TL;DR Anthropic-Helpful
Method T = 0.001 T = 0.5 T = 1.0 Average T = 0.001 T = 0.5 T = 1.0 Average
SFT 46.6 44.9 39.8 43.8 59.1 56.2 56.8 57.4
DPO Rafailov et al. (2024) 66.5 67.0 69.5 67.7 73.5 72 69.5 71.7
KTO Ethayarajh et al. (2024) 71.8 71.9 70.6 71.4 72.8 72.9 68.8 71.5
JPO (Ours) 72.7 71.9 74.2 72.9 76.3 74.5 74.1 75.0
Table 5: Results for aligning LLMs with the JPO preference optimization objective. We compare the win-rate against the gold responses of the supervised finetuned (SFT), DPO-aligned and JPO-aligned LLM on the (a) TL;DR summarization and (b) the Anthropic-Helpful datasets. In our experiments, we utilize ChatGPT to compare the model responses with the gold responses. We generate model responses for three sampling temperatures. The results are averaged over three runs of the preference optimization objectives.
Appendix F Qualitative Examples
@@
F.1 Anthropic-Helpful Examples
We present the qualitative examples for the preferences acquired for the Anthropic-helpful dataset in Figure 8, and 9. We present our observations in the figure captions.
@@ F.1 Anthropic-Helpful Examples
F.2 TL;DR Summarization Examples
We present the qualitative examples for the preferences acquired for the TL;DR summarization dataset in Figure 10, 11, and 12. We present our observations in the figure captions.
@@ F.2 TL;DR Summarization Examples
Appendix G Alignment Training Details
G.1 Supervised Finetuning Details
We present the SFT details in table  6. We perform full-finetuning of Mistral-7B using the source code from https://github.com/abacaj/fine-tune-mistral.
@@ G.1 Supervised Finetuning Details
G.2 JPO
We present the training details for JPO preference optimization objective in the Table 7. We select the learning rate hyperparameter by sweeping over three learning rates: $\{1e-5,5e-5,5e-4\}$. We utilize the TRL library von Werra et al. (2020) for the DPO source code.
@@ G.2 JPO
Optimizer AdamW (Loshchilov and Hutter, 2017)
@@
Table 7: Training details for JPO preference optimization objective. We use the identical settings for DPO.
Appendix H ChatGPT Prompts
You are a helpful following assistant whose goal is to select the preferred output for a given instruction. Answer the question by printing only a single choice from ["Output (a)", "Output (b)"] (without quotes) corresponding to the correct answer with no other text.
#### Annotation Guideline
You will read examples, which are composed of the following:
1. an Instruction we give to the AI system
2. Output (a), the first output from the AI system
3. Output (b), the second output from the AI system
### Instruction:
${instruction}$
### Output (a):
${output_{1}}$
### Output (b):
${output_{2}}$
## Preferred Output:
You are a helpful following assistant whose goal is to select the preferred output for a given instruction. Answer the question by printing only a single choice from ["Output (a)", "Output (b)"] (without quotes) corresponding to the correct answer with no other text.
#### Annotation Guideline
You will read examples, which are composed of the following:
### Instruction (a):
${instruction_{1}}$
### Output (a):
${output_{1}}$
### Instruction (b):
${instruction_{2}}$
### Output (b):
${output_{2}}$
## Preferred Output:
Appendix I Human Annotation Platform
For human evaluation, we recruit annotators from Amazon Mechanical Turk, and all annotators are fairly paid more than \$18 USD per hour (it varies depending on the time spent on HITs), which is higher than the national minimum wage where the annotators are recruited. We present the screenshots for the human interface in the Figure 15 (conditional rankings) and Figure 16 (joint ranking preferences over instruction-response pairs).
Figure 15: Human annotation interface for Conditional Rankings